# NCL-UoR at SemEval-2025 Task 3: Detecting Multilingual Hallucination and Related Observable Overgeneration Text Spans with Modified RefChecker and Modified SeflCheckGPT

### v1

@@ Abstract
SemEval-2025 Task 3 (Mu-SHROOM) focuses on detecting hallucinations in content generated by various large language models (LLMs) across multiple languages. This task involves not only identifying the presence of hallucinations but also pinpointing their specific occurrences. To tackle this challenge, this study introduces two methods: modified RefChecker and modified SelfCheckGPT. The modified RefChecker integrates prompt-based factual verification into References, structuring them as claim-based tests rather than single external knowledge sources. The modified SelfCheckGPT  incorporates external knowledge to overcome its reliance on internal knowledge. In addition, both methods’ original prompt designs are enhanced to identify hallucinated words within LLM-generated texts. Experimental results demonstrate the effectiveness of the approach, achieving a high ranking on the test dataset in detecting hallucinations across various languages, with an average IoU of 0.5310 and an average COR of 0.5669. The source code used in this paper is available at https://github.com/jianfeixu95/NCL-UoR.
@@
Jiaying Hong1, Thanet Markchom2, Jianfei Xu1, Tong Wu3and Huizhi Liang1 1 School of Computing, Newcastle University, Newcastle upon Tyne, UK 2 Department of Computer Science, University of Reading, Reading, UK 3 Formerly at School of Computing, Newcastle University, Newcastle upon Tyne, UK hongjalynn@gmail.com, thanet.markchom@reading.ac.uk, mr.xujianfei@gmail.com, tongwuwhitney@gmail.com, huizhi.liang@newcastle.ac.uk
@@ 2.1 Modified-RefChecker (MRC)
MRC  is an improved RefChecker, integrating CLAUDE Anthropic (2022) for enhanced functionality. It consists of two key components: the Extractor for constructing references and the Checker for identifying hallucinated words along with their probabilities. Figure 1 shows the overview of MRC. The details of each component are described below.
@@ 2.2 Modified-SelfCheckGPT-H (MSCGH)
To detect hallucination words, an LLM (this work considers GPT-3.5, GPT-4 and GPT-4o) is used to answer the prompt created in the previous step for each LLM output text. For each response, the hallucination words are identified, and a list of index intervals indicating the positions of these words in the LLM output string, =$\mathcal{L}=\{L_{1},L_{2},\dots,L_{n}\}$, is obtained. Then, all overlapping and adjacent intervals across all $N$ responses are merged into a set of distinct, non-overlapping intervals =$\mathcal{M}=\{(s_{1},e_{1}),(s_{2},e_{2}),\dots,(s_{m},e_{m})\}$.
@@ 4 Results and Discussions
Across multiple languages, the top three models exhibited distinct performance differences, as shown in Figure 3. MSCGH_GPT_CSE_F  consistently led, benefiting from more effective keyword extraction, comprehensive external knowledge retrieval, and stronger hallucination detection. Its broader retrieval strategy provided an advantage in handling ambiguous or multi-step queries. MSCGH_GPT_WIKI_A  followed closely, particularly excelling in Chinese results, where complex segmentation and word relationships were better handled through its customized keyword extraction. MRC_CLAUDE_CSE_A, while still effective, showed greater variance across languages, likely due to less optimized retrieval strategies or weaker hallucination detection.
@@
Figure 4 compares different LLMs. From this figure, CLAUDE performed the worst, while GPT-4o showed significant improvements. However, not all GPT-4o-based methods outperformed CLAUDE, indicating that LLM upgrades alone do not guarantee better results—effective knowledge retrieval remained essential. Performance gaps between LLMs were more pronounced in high-resource languages (e.g., Italian), where GPT-4o significantly outperformed CLAUDE. In contrast, for low-resource languages (e.g., Arabic), GPT-4o’s benefits were inconsistent—some methods showed only marginal gains, while those using the Google CSE API achieved substantial improvements. This underscored the critical role of external knowledge integration in maximizing LLM performance.
@@ 5 Conclusion
SemEval-2025 Mu-SHROOM introduced the task of detecting hallucination spans in multilingual LLM outputs. To tackle this task, this work proposed two methods: modified RefChecker (MRC) and modified SelfCheckGPT (MSCGH). These methods incorporated external knowledge integration and an improved prompt design, enabling the detection of text-span hallucinations in LLM-generated texts. MRC and variations of MSCGH  (with different keyword extraction techniques, external knowledge sources, and prompt strategies) were evaluated across datasets in 14 languages. Three top-performing methods were chosen for discussion in this paper. Among the evaluated methods, MSCGH using GPT-3.5 for keyword extraction, full Google CSE API results, and GPT-4o for hallucination detection achieved the best overall performance. Although MSCGH  demonstrated higher performance, it lacked stability when applied across different languages. Meanwhile, MRC was more stable but less optimized. Future research could focus on refining the prompt design and enhancing external knowledge integration strategies. Additionally, adaptive learning for low-resource languages and broader language task expansion could be considered.

### v2

@@ Abstract
SemEval-2025 Task 3 (Mu-SHROOM) focuses on detecting hallucinations in content generated by various large language models (LLMs) across multiple languages. This task involves not only identifying the presence of hallucinations but also pinpointing their specific occurrences. To tackle this challenge, this study introduces two methods: Modified-RefChecker (MRC) and Modified-SelfCheckGPT-H (MSCGH). MRC integrates prompt-based factual verification into References, structuring them as claim-based tests rather than single external knowledge sources. MSCGH incorporates external knowledge to overcome its reliance on internal knowledge. In addition, both methods’ original prompt designs are enhanced to identify hallucinated words within LLM-generated texts. Experimental results demonstrate the effectiveness of the approach, achieving a high ranking on the test dataset in detecting hallucinations across various languages, with an average IoU of 0.5310 and an average COR of 0.5669. The source code used in this paper is available at https://github.com/jianfeixu95/NCL-UoR.
@@
Jiaying Hong1, Thanet Markchom2, Jianfei Xu1, Tong Wu3and Huizhi Liang1 1 School of Computing, Newcastle University, Newcastle upon Tyne, UK 2 Department of Computer Science, University of Reading, Reading, UK 3 Previously at School of Computing, Newcastle University, Newcastle upon Tyne, UK hongjalynn@gmail.com, thanet.markchom@reading.ac.uk, mr.xujianfei@gmail.com, tongwuwhitney@gmail.com, huizhi.liang@newcastle.ac.uk
@@ 1 Introduction
Overall, unlike the original RefChecker and SelfCheckGPT, which rely on static references and internal prompt-based self-consistency, respectively, our modified methods incorporate external knowledge retrieval and prompt-driven span-level verification to improve hallucination detection accuracy and granularity.
2 Related Work
Most recent approaches to detecting hallucinations in LLM outputs rely on prompting techniques, where the models evaluate the likelihood of hallucinations in their responses. For instance, Kadavath et al. (2022) proposed prompting LLMs to generate an answer and then predict the probability of its correctness. Manakul et al. (2023) introduced SelfCheckGPT, which compares an LLM-generated sentence against multiple alternative generations, asking the model to assess whether the original sentence is consistently supported. Friel and Sanyal (2023) presented ChainPoll, using detailed prompts to guide models in identifying hallucinations. Hu et al. (2024) proposed RefChecker, a retrieval-augmented evaluation method that checks the consistency of model outputs against retrieved external references, aiming to identify factual inconsistencies and hallucinations without relying solely on LLM self-judgment. However, most existing methods focus on detecting whether a text contains hallucinations or not. Identifying the specific parts of a text that are hallucinations remains an open research challenge. Therefore, in this work, we modified RefChecker and SelfCheckGPT, two state-of-the-art methods to handle this task.
@@ 3.1 Modified-RefChecker (MRC)
MRC  is an improved RefChecker, integrating CLAUDE Anthropic (2022) for enhanced functionality. Note that any LLM, including open-source ones, can be substituted. However, to ensure consistent and scalable evaluation, we adopt CLAUDE due to its multilingual support, API stability, and superior performance compared to open-source models in the original RefChecker Hu et al. (2024). MRC consists of two key components: the Extractor for constructing references and the Checker for identifying hallucinated words along with their probabilities. Figure 1 shows the overview of MRC. The details of each component are described below.
@@ 3.2 Modified-SelfCheckGPT-H (MSCGH)
To detect hallucination words, an LLM (this work considers GPT-3.5, GPT-4 and GPT-4o following the original methodology of using GPT models in SelfCheckGPT Manakul et al. (2023).) is used to answer the prompt created in the previous step for each LLM output text. For each response, the hallucination words are identified, and a list of index intervals indicating the positions of these words in the LLM output string, =$\mathcal{L}=\{L_{1},L_{2},\dots,L_{n}\}$, is obtained. Then, all overlapping and adjacent intervals across all $N$ responses are merged into a set of distinct, non-overlapping intervals =$\mathcal{M}=\{(s_{1},e_{1}),(s_{2},e_{2}),\dots,(s_{m},e_{m})\}$.
@@ 4 Datasets and Experimental Setup
Baseline
Three baselines were provided in the task Vázquez et al. (2025): (1) Baseline (neural): Fine-tuning of the neural network classifier based on XLM-R, outputting binary (0/1) probability predictions for each token, (2) Baseline (mark-all): Predicting all characters as hallucinations ($probability=1$), and (3) Baseline (mark-none): Predicting all characters as non-hallucinations ($probability=0$).
@@ 5 Results and Discussions
The comparative results of our methods and the baselines are shown in Figures  3(a) and  3(b). From these figures, the neural and mark-none baselines performed the worst across all languages, while the mark-all baseline achieved slightly higher IoU but nearly zero COR scores. In contrast, our method outperformed these baselines in all languages, with average improvements of approximately 0.30 in IoU and 0.45 in COR. More importantly, according to the 100,000 bootstrap resamplings mentioned in Vázquez et al. (2025), our submitted methods achieved a $Pr(rank)$ above 0.5 in every language. This indicates a higher probability of outperforming the next-best team in the majority of samples, and thus demonstrates robust and consistent cross-lingual performance.
Across multiple languages, our models exhibited distinct performance differences, as shown in Figure 3. MSCGH_GPT_CSE_F  consistently led, benefiting from more effective keyword extraction, comprehensive external knowledge retrieval, and stronger hallucination detection. Its broader retrieval strategy provided an advantage in handling ambiguous or multi-step queries. MSCGH_GPT_WIKI_A  followed closely, particularly excelling in Chinese results, where complex segmentation and word relationships were better handled through its customized keyword extraction. MRC_CLAUDE_CSE_A, while still effective, showed greater variance across languages, likely due to less optimized retrieval strategies or weaker hallucination detection.
@@
Figure 4 compares different LLMs. In this figure, each model is represented with bars showing the IoU and COR scores for individual languages. The grey bar behind each model’s score bars indicates the average IoU and COR scores for that model. From this figure, CLAUDE performed the worst, while GPT-4o showed significant improvements. However, not all GPT-4o-based methods outperformed CLAUDE, indicating that LLM upgrades alone do not guarantee better results—effective knowledge retrieval remained essential. Performance gaps between LLMs were more pronounced in high-resource languages (e.g., Italian), where GPT-4o significantly outperformed CLAUDE. In contrast, for low-resource languages (e.g., Arabic), GPT-4o’s benefits were inconsistent—some methods showed only marginal gains, while those using the Google CSE API achieved substantial improvements. This underscored the critical role of external knowledge integration in maximizing LLM performance.
@@ 6 Conclusion
SemEval-2025 Mu-SHROOM introduced the task of detecting hallucination spans in multilingual LLM outputs. To tackle this task, this work proposed two methods: Modified-RefChecker (MRC) and Modified-SelfCheckGPT-H (MSCGH). These methods incorporated external knowledge integration and an improved prompt design, enabling the detection of text-span hallucinations in LLM-generated texts. MRC and variations of MSCGH  (with different keyword extraction techniques, external knowledge sources, and prompt strategies) were evaluated across datasets in 14 languages. Three top-performing methods were chosen for discussion in this paper. Among the evaluated methods, MSCGH using GPT-3.5 for keyword extraction, full Google CSE API results, and GPT-4o for hallucination detection achieved the best overall performance. Although MSCGH  demonstrated higher performance, it lacked stability when applied across different languages. Meanwhile, MRC was more stable but less optimized. One limitation of the proposed approaches is the assumption that external knowledge is accurate. However, the retrieved information may not always be fully factual due to the ever-growing volume of online content. Such inaccuracies could reduce the effectiveness of the proposed approaches. Future research could focus on refining the prompt design and enhancing external knowledge integration and faulty correction strategies. Additionally, adaptive learning for low-resource languages and broader language task expansion could be considered.